# Point-In-Context: Understanding Point Cloud via In-Context Learning

### v1

@@ 4.2 Point-In-Context-Segmenter Model
Although PIC-G successfully demonstrates multitasking capabilities through the MPM framework with the JS module, it still has limitations in segmentation tasks. Therefore, to explore a PIC expert in the field of segmentation, we further propose another version of PIC called Point-In-Context-Segmenter (PIC-S), comprising two primary novel training strategies: In-Context Labeling and In-Context Enhancing.
@@
Limitations of PIC-G on Segmentation Tasks. In the PIC-G framework, to adjust to the input-output formats of other point cloud tasks, all part labels in the segmentation task are converted into points with XYZ coordinates in space, forming a pre-defined label point map. The label point map is fixed during training, leading to two main issues: First, fixed label points make the model rely solely on the static three-dimensional positions of label points for point cloud segmentation. This is not conducive to learning the mapping relationship in example pairs (prompts), contradicting the original intention of in-context learning. Therefore, part segmentation cannot gain mutual benefit from other sub-tasks. Second, when additional part segmentation datasets are incorporated, the large set of label points crowd in the limited space, exacerbating the model’s difficulty in regressing prediction to the correct partition. As a result, a small regression error has a higher possibility of leading to a wrong prediction. Third, once the label-point assignment is fixed, the space partition is determined that each label covers a specific region. In this case, allocating new space to novel classes is challenging. Therefore, the fixed label choice sacrifices the possibility of generalization.
In-Context Labeling. To address the shortcomings of PIC-G discussed above, we propose In-Context Labeling. Instead of using the pre-defined label-point assignment, we employ dynamic in-context labels. Specifically, the contextual labeling process involves three steps (also illustrated in Fig. 4).
Figure 4: Comparison of generating targets between In-Context Labeling (PIC-S) and pre-defined label map (PIC-G). $P$ represents the number of all parts ($P\gg N_{B}>|C_{i}|$).
Firstly, we define the label points bank $B=\{p_{k}\in\mathbb{R}^{3},k=1,2,\cdots,N_{B}\}$ of size $N_{B}$ by evenly split the 3D coordinate space, so that
$N_{B}\geq\underset{i\leq n}{\max}|C_{i}|,$ (4)
where $C_{i}$ is the label set of $i$-th sample and $n$ is the dataset size.
The second step involves re-labeling each point cloud in the segmentation task. Unlike the previous way where one fixed label point represents one component, we randomly assign label points from $B$ to the parts within that point cloud each time we take a point cloud from the dataset. Specifically, we randomly select $|C_{i}|$ label points from the pre-defined label point bank to form the label point set $B_{i}$ of the $i$-th point cloud:
$B_{i}=\{p_{k}\in\mathbb{R}^{3},k=1,2,...,|C_{i}|\}\in B.$ (5)
Then we allocate them randomly to the parts contained within the point cloud. With such random assignment, the binding of labels to label points is broken, thus allowing for novel categories.
The third step is to select a prompt $Q_{j}$ that shares the same part semantics so as the label set $B_{i}$. This is essential as prompts play a crucial role in in-context learning. The more similar are $<Q_{i},Q_{j}>$ and $<T_{i},T_{j}>$, the more guidance can be injected into the in-context learning model on the fly.
@@
Our context-based labeling approach allows the model to focus more on the semantic information presented in the prompts. Besides, it enables our framework to easily incorporate additional segmentation datasets without increasing the complexity of identifying point cloud semantic information. During training, the model also segments the query point cloud by observing the hidden pattern between the input point cloud and the label point in the prompt rather than relying on the label point’s three-dimensional coordinates. Thus, the in-context labeling provides the scalability and extensibility for the model to generalize to other segmentation datasets. We only need to randomly assign label points to the point clouds in out-of-domain segmentation datasets and provide a prompt that shares the same label points.
In-Context Enhancing. As discussed above, we introduce a new training method of in-context labeling that involves completely randomly assigning labels to label points. However, such randomness presents a certain challenge for the model training. To further stimulate the in-context learning capability of the model, we introduce the In-Context Enhancing training strategy. Specifically, we first apply random corruption operations to a clean point cloud $P_{j}$ and construct pairs of point clouds aiming to restore the damaged point cloud, so the query point cloud pair can be formulated as:
$Q^{r}_{j}=\{r(P_{j}),P_{j}\},$ (6)
where $r(\cdotp)$ is corruption functions, including common point cloud corruption operations. Each point cloud $P_{j}$ only contains XYZ coordinates without any color or component semantic information.
Additionally, we provide a pair of point clouds that have undergone the same corruption operation $r(\cdotp)$ as a prompt, requiring the model to restore the damaged point cloud based on the provided prompt. Subsequently, we involve the constructed pairs of point clouds in the training set to enhance the model’s in-context learning capability during training. This approach not only diversifies the training data but also encourages the model to develop robust restoration techniques, thereby improving its overall performance in handling real-world scenarios.
Furthermore, by focusing solely on the XYZ coordinates of each point cloud without any additional color or semantic information, we ensure that the model learns to restore point clouds based on spatial context alone. This abstraction fosters a deeper understanding of the underlying structure of point clouds, enabling the model to learn more effectively within segmentation point cloud pairs across diverse datasets, which can be proved in Sec. 5.3.
@@
TABLE I: Comparison results on ShapeNet In-Context Datasets, including reconstruction, denoising, registration, and part segmentation. PIC-G and PIC-S represent Point-In-Context-Generalist (Sec.4.1) and Point-In-Context-Segmenter (Sec.4.2), respectively. We also report task-specific and multitask models. $\dagger$ indicates models that consist of a shared backbone and multiple task heads. $\ddagger$ indicates models whose backbone is pre-trained via specific tasks. For reconstruction, denoising, and registration, we report the Chamfer Distance [104] (CD) loss (x1000). For part segmentation, we report the mean Intersection over Union (mIoU) metric. Bold: best. Underline: second-best.
@@
PointNet [38] CVPR’17 3.7 3.7 3.8 3.9 4.1 3.9 4.1 4.0 4.1 4.0 4.2 4.1 5.3 5.9 6.9 7.7 8.5 6.9 77.45
DGCNN [41] TOG’19 3.9 3.9 4.0 4.1 4.3 4.0 4.7 4.5 4.6 4.5 4.7 4.6 6.2 6.7 7.3 7.4 7.7 7.1 76.12
PCT [48] CVM’21 2.4 2.4 2.5 2.6 3.0 2.6 2.3 2.2 2.2 2.2 2.3 2.2 5.3 5.7 6.3 6.9 7.2 6.3 79.46
ACT [74] ICLR’23 2.4 2.5 2.3 2.5 2.8 2.5 2.2 2.3 2.2 2.3 2.5 2.3 5.1 5.6 5.9 6.0 7.0 5.9 81.24
@@
PointNet [38] CVPR’17 87.2 86.6 87.3 90.8 92.2 88.8 17.8 22.0 25.6 30.4 33.2 25.8 25.4 22.6 24.9 25.7 26.9 25.1 15.33
DGCNN [41] TOG’19 38.8 36.6 37.5 37.9 42.9 37.7 6.5 6.3 6.5 6.4 7.1 6.5 12.5 14.9 17.9 19.7 20.7 17.1 16.95
PCT [48] CVM’21 34.7 44.1 49.9 50.0 52.3 46.2 11.2 10.3 10.7 10.2 10.5 10.6 24.4 26.0 29.6 32.8 34.7 29.5 16.71
@@
PointNet [38] CVPR’17 47.0 45.8 45.4 45.4 45.8 45.9 22.9 23.2 26.3 28.3 30.0 26.1 35.5 34.8 37.1 37.2 38.6 36.6 10.13
DGCNN [41] TOG’19 46.7 47.2 48.1 48.6 48.5 47.8 8.2 8.3 8.4 8.8 9.2 8.6 14.2 15.8 18.2 21.8 23.5 18.7 21.35
PCT [48] CVM’21 64.7 60.8 59.2 60.1 59.7 61.0 14.5 12.2 12.4 12.0 11.8 12.6 22.6 25.2 28.3 31.1 33.2 28.1 15.43
PointMAE [27] ECCV’22 5.5 5.5 6.1 6.4 6.4 6.0 5.6 5.4 5.6 5.5 5.8 5.6 11.4 12.8 14.8 16.0 16.9 14.5 5.42
ACT [74] ICLR’23 7.4 6.6 6.5 6.6 7.0 6.8 7.3 6.8 7.0 6.8 7.2 7.0 12.2 14.4 19.4 25.5 29.0 20.1 12.08
I2P-MAE [28] CVPR’23 17.0 16.0 16.7 17.2 18.5 17.2 20.6 20.4 20.1 18.3 18.8 19.6 32.5 31.3 31.1 31.6 31.2 31.5 22.60
ReCon [73] ICML’23 12.4 12.1 12.4 12.5 13.1 12.5 20.4 24.5 27.2 29.2 32.5 26.9 14.7 16.3 19.2 21.5 22.5 18.8 7.71
@@
Point-BERT [67] CVPR’22 288 285 292 286 308 292 292 293 298 296 299 296 291 295 294 295 298 294 0.65
Our PIC-G-Cat / 3.2 3.6 4.6 4.9 5.5 4.3 3.9 4.6 5.3 6.0 6.8 5.3 10.0 11.4 13.8 16.9 18.6 14.1 78.95
Our PIC-G-Sep / 4.7 4.3 4.3 4.4 5.7 4.7 6.3 7.2 7.9 8.2 8.6 7.6 8.6 9.2 10.2 11.3 12.4 10.3 74.95
Our PIC-S-Cat / 9.3 5.1 4.8 5.0 10.3 6.9 4.7 5.7 6.5 7.4 8.2 6.5 12.8 15.8 23.9 31.2 36.9 24.1 83.78
Our PIC-S-Sep / 4.6 4.5 4.5 4.8 7.1 5.1 9.4 11.7 12.5 13.1 13.4 12.0 6.0 6.1 7.6 6.7 7.3 6.7 83.73

### v2

@@ 4.2 Point-In-Context++ Model
Although PIC successfully demonstrates multitasking capabilities through the MPM framework with the JS module, it still has limitations in segmentation tasks. Therefore, to empower the segmentation capability of PIC, we further propose another version of PIC called PIC++, comprising two primary novel training strategies: In-Context Labeling and In-Context Enhancing.
Algorithm 1 In-Context Labeling: Construction and Dynamic Assignment
1:Label Bank Size $N_{B}$, Point Cloud Dataset $\mathcal{D}$
2:Label Point Bank $B$, Mapped Labels for Batch
3:Stage 1: Constructing the Label Point Bank
4:Initialize empty bank $B\leftarrow\emptyset$
5:for $k=1$ to $N_{B}$ do
6:  Sample $p_{k}\sim\text{Uniform}([0,1]^{3})$ // Sample from unit cube
7:  $B\leftarrow B\cup\{p_{k}\}$
8:end for
9:
10:Stage 2: Dynamic Re-labeling (Per Iteration)
11:while Training or Inference do
12:  Get current object category $C$ and its semantic parts list $L_{C}=\{l_{1},l_{2},\dots,l_{m}\}$
13:  // Randomly select subset from Bank
14:  Sample indices $I_{sub}\subset\{1,\dots,N_{B}\}$ such that $|I_{sub}|=m$
15:  Extract temporary label points $B_{sub}=\{B[i]\mid i\in I_{sub}\}$
16:  // Establish Mapping: Semantic Part $\to$ Coordinate
17:  Initialize Mapping $\Phi\leftarrow\emptyset$
18:  for $j=1$ to $m$ do
19:   $\Phi(l_{j})\leftarrow B_{sub}[j]$
20:  end for
21:  // Apply Consistent Mapping to Prompt and Query
22:  Prompt Labels $Y_{prompt}\leftarrow\{\Phi(y)\mid y\in\text{Prompt Part Categories}\}$
23:  Query Labels $Y_{query}\leftarrow\{\Phi(y)\mid y\in\text{Query Part Categories}\}$
24:  Proceed with Model Optimization/Inference using $Y_{prompt}$ and $Y_{query}$
25:end while
@@
Limitations of PIC on Segmentation Tasks. In the PIC framework, to adjust to the input-output formats of other point cloud tasks, all part labels in the segmentation task are converted into points with XYZ coordinates in space, forming a pre-defined label point map. The label point map is fixed during training, leading to two main issues: First, fixed label points make the model rely solely on the static three-dimensional positions of label points for point cloud segmentation. This is not conducive to learning the mapping relationship in example pairs (prompts), contradicting the original intention of in-context learning. Therefore, part segmentation cannot gain mutual benefit from other sub-tasks. Second, when additional part segmentation datasets are incorporated, the large set of label points crowd in the limited space, exacerbating the model’s difficulty in regressing prediction to the correct partition. As a result, a small regression error has a higher possibility of leading to a wrong prediction. Third, once the label-point assignment is fixed, the space partition is determined that each label covers a specific region. In this case, allocating new space to novel classes is challenging. Therefore, the fixed label choice sacrifices the possibility of generalization.
In-Context Labeling. To address the limitations of the static assignment in PIC, we propose a dynamic strategy named In-Context Labeling (ICL). Unlike previous methods that rely on a fixed one-to-one binding between label points and semantic classes, ICL employs a dynamic, context-aware assignment mechanism. The process consists of three key steps, as illustrated in Fig. 4 and Algorithm. 1:
Constructing the Label Point Bank. First, we initialize a comprehensive label point bank $B=\{p_{k}\in\mathbb{R}^{3},k=1,2,\cdots,N_{B}\}$ by uniformly sampling the 3D coordinate space. The size of the bank, $N_{B}$, is set to satisfy:
$N_{B}\geq\max_{i\leq n}|C_{i}|,$ (4)
where $C_{i}$ denotes the number of semantic categories in the $i$-th sample, and $n$ represents the total dataset size.
Dynamic Re-labeling. The second step involves dynamically assigning labels for each training iteration. In contrast to traditional methods where a specific spatial point permanently represents a specific component (e.g., a fixed point for “wing”), we break this static binding. For each input point cloud $i$, we randomly sample $|C_{i}|$ points from the bank $B$ to construct a temporary label set $B_{i}$:
$B_{i}\subset B,\quad\text{s.t.}\quad|B_{i}|=|C_{i}|.$ (5)
These selected points are then randomly mapped to the semantic parts of the current point cloud. This stochastic assignment forces the model to decouple semantic meaning from absolute spatial coordinates, thereby facilitating the learning of novel categories.
Prompt Consistency. Finally, to enable in-context learning, we select a prompt $R_{i}$ (support sample) that shares the same semantic-to-label mapping as the query input $Q_{j}$ from the training set. This consistency is crucial; it serves as a contextual anchor. By observing how the label points in $B_{i}$ are applied to the prompt $R_{i}$, the model infers the segmentation rule for the current query.
@@
Benefits. Our context-based labeling approach allows the model to focus more on the semantic information presented in the prompts. Besides, it enables our framework to easily incorporate additional segmentation datasets without increasing the complexity of identifying point cloud semantic information. During training, the model also segments the query point cloud by observing the hidden pattern between the input point cloud and the label point in the prompt rather than relying on the label point’s three-dimensional coordinates. Thus, the in-context labeling provides the scalability and extensibility for the model to generalize to other segmentation datasets. We only need to randomly assign label points to the point clouds in out-of-domain segmentation datasets and select a random sample from the same category as the prompt. This ensures the prompt and query share consistent label definitions, a process that is fully automated without handcrafted intervention.
@@
Table 2: Comparison results on ShapeNet In-Context Datasets, including reconstruction, denoising, registration, and part segmentation. PIC and PIC++ represent models described in Sec.4.1 and Sec.4.2 respectively. Our PICs only need to be trained once and do not require fine-tuning or any parameter updates during testing. We also report task-specific and multitask models. $\dagger$ indicates models with a shared backbone and multiple task heads. $\ddagger$ indicates models whose backbone is pre-trained via specific tasks. For reconstruction, denoising, and registration, we report the Chamfer Distance [chamferdistance] (CD) loss ($\times$1000). We report the mean Intersection over Union (mIoU) metric for part segmentation. Bold: best. Underline: second-best. red and blue arrows indicate whether PIC++ is worse or better than PIC respectively.
@@
PointNet [pointnet]  CVPR’17    3.7 3.7 3.8 3.9 4.1    3.9    4.1 4.0 4.1 4.0 4.2    4.1    5.3 5.9 6.9 7.7 8.5    6.9    77.45
DGCNN [dgcnn]  TOG’19    3.9 3.9 4.0 4.1 4.3    4.0    4.7 4.5 4.6 4.5 4.7    4.6    6.2 6.7 7.3 7.4 7.7    7.1    76.12
PCT [pct]  CVM’21    2.4 2.4 2.5 2.6 3.0    2.6    2.3 2.2 2.2 2.2 2.3    2.2    5.3 5.7 6.3 6.9 7.2    6.3    79.46
ACT [act]  ICLR’23    2.4 2.5 2.3 2.5 2.8    2.5    2.2 2.3 2.2 2.3 2.5    2.3    5.1 5.6 5.9 6.0 7.0    5.9    81.24
@@
PointNet [pointnet]  CVPR’17    87.2 86.6 87.3 90.8 92.2    88.8    17.8 22.0 25.6 30.4 33.2    25.8    25.4 22.6 24.9 25.7 26.9    25.1    15.33
DGCNN [dgcnn]  TOG’19    38.8 36.6 37.5 37.9 42.9    37.7    6.5 6.3 6.5 6.4 7.1    6.5    12.5 14.9 17.9 19.7 20.7    17.1    16.95
PCT [pct]  CVM’21    34.7 44.1 49.9 50.0 52.3    46.2    11.2 10.3 10.7 10.2 10.5    10.6    24.4 26.0 29.6 32.8 34.7    29.5    16.71
@@
PointNet [pointnet]  CVPR’17    47.0 45.8 45.4 45.4 45.8    45.9    22.9 23.2 26.3 28.3 30.0    26.1    35.5 34.8 37.1 37.2 38.6    36.6    10.13
DGCNN [dgcnn]  TOG’19    46.7 47.2 48.1 48.6 48.5    47.8    8.2 8.3 8.4 8.8 9.2    8.6    14.2 15.8 18.2 21.8 23.5    18.7    21.35
PCT [pct]  CVM’21    64.7 60.8 59.2 60.1 59.7    61.0    14.5 12.2 12.4 12.0 11.8    12.6    22.6 25.2 28.3 31.1 33.2    28.1    15.43
PointMAE [pointmae]  ECCV’22    5.5 5.5 6.1 6.4 6.4    6.0    5.6 5.4 5.6 5.5 5.8    5.6    11.4 12.8 14.8 16.0 16.9    14.5    5.42
ACT [act]  ICLR’23    7.4 6.6 6.5 6.6 7.0    6.8    7.3 6.8 7.0 6.8 7.2    7.0    12.2 14.4 19.4 25.5 29.0    20.1    12.08
I2P-MAE [i2pmae]  CVPR’23    17.0 16.0 16.7 17.2 18.5    17.2    20.6 20.4 20.1 18.3 18.8    19.6    32.5 31.3 31.1 31.6 31.2    31.5    22.60
ReCon [recon]  ICML’23    12.4 12.1 12.4 12.5 13.1    12.5    20.4 24.5 27.2 29.2 32.5    26.9    14.7 16.3 19.2 21.5 22.5    18.8    7.71
MAMBA3D [han2024mamba3d]  MM’24    8.4 7.8 9.3 9.8 10.5    9.2    13.2 13.8 13.5 14.8 12.8    13.6    9.9 13.2 13.9 19.7 20.3    15.4    14.58
PCMamba [pcm]  AAAI’25    8.1 8.9 10.2 10.5 11.7    9.9    14.5 14.2 15.9 14.7 14.5    14.8    11.1 13.8 18.2 21.8 22.7    17.5    19.53
@@
Point-BERT [pointbert]  CVPR’22    288 285 292 286 308    292    292 293 298 296 299    296    291 295 294 295 298    294    0.65
”-Cat” PIC  NeurIPS’23    3.2 3.6 4.6 4.9 5.5    4.3  3.9 4.6 5.3 6.0 6.8    5.3  10.0 11.4 13.8 16.9 18.6    14.1    78.95
PIC++  /    4.5 3.3 3.7 4.6 5.4  4.3-  3.8 4.3 5.1 5.6 6.7  5.1  9.7 11.6 12.8 15.9 18.0  13.6  85.27
”-Sep” PIC  NeurIPS’23    4.7 4.3 4.3 4.4 5.7    4.7    6.3 7.2 7.9 8.2 8.6    7.6    8.6 9.2 10.2 11.3 12.4    10.3  74.95
PIC++  /    4.5 4.1 4.2 4.2 5.9  4.6  5.8 7.1 7.7 8.2- 8.4  7.4  7.6 6.2 6.8 6.9 7.9  7.1  85.53
In-Context Enhancing. As discussed above, we introduce a new training method of in-context labeling that involves completely randomly assigning labels to label points. This decoupling of semantics from fixed coordinates prevents the model from relying on rote memorization, but it also imposes a higher requirement on the model’s ability to infer mapping rules from the context. To bridge this gap and further stimulate the in-context learning capability of the model, we introduce the In-Context Enhancing (ICE) training strategy.
Unlike static multi-task benchmarks (e.g., ShapeNet In-Context) that use fixed pre-defined tasks, we first dynamically apply random corruption operations to a clean point cloud $P_{j}$ and construct pairs of point clouds aiming to restore the damaged point cloud. The query point cloud pair can be formulated as:
$Q^{r}_{j}=\{r(P_{j}),P_{j}\},$ (6)
where $r(\cdotp)$ is a stochastic corruption function encompassing a diverse range of point cloud corruption operations (e.g., local masking, jittering, dropping). Each point cloud $P_{j}$ only contains XYZ coordinates without any color or component semantic information.
Additionally, we provide a pair of point clouds that have undergone the same corruption operation $r(\cdotp)$ as a prompt, requiring the model to restore the damaged point cloud based on the provided prompt. Subsequently, we involve the constructed pairs of point clouds in the training set to enhance the model’s in-context learning capability during unified training (as illustrated in Fig. 4 (b)).
Critically, ICE serves as a geometric pretext task designed to synergize with In-Context Labeling. By forcing the model to identify varying degradation patterns in the prompt and transfer the restoration logic to the query, we foster a generalized “meta-capability” for analyzing prompt-query relationships. This differs fundamentally from standard multi-task learning, as the goal is not merely to solve the restoration task, but to train the model’s sensitivity to contextual rules. Furthermore, by focusing solely on the XYZ coordinates of each point cloud without any additional color or semantic information, we ensure that the model learns to restore point clouds based on spatial context alone. This abstraction fosters a deeper understanding of the underlying structure of point clouds, enabling the model to learn more effectively within segmentation pairs across diverse datasets, which can be proved in Sec. 5.3.
Figure 5: Visualization of predictions from the ”-Sep” variant of PIC on ShapeNet In-Context Datasets. For part segmentation, we visualize the generated target together with the mapping back, both adding category-specific colors for a better look.
Figure 6: Visualization of comparison results between PIC and multitask models on reconstruction (lines 1-2), denoising (lines 3-4), and registration (lines 5-6), where our models can generate more accurate predictions than other multitask models.
Figure 7: (a) Generalized to out-of-distribution data. We evaluate the registration task on the ModelNet40 [modelnet40] dataset, which is not present in the training set. The Chamfer Distance (x1000) is used as the evaluation metric. (b) Generalized to new tasks. Our model can rotate at any angle according to the example pairs and show the completion ability to the broken point cloud, which can be treated as an upgraded task of registration.
Inference of PIC++. In contrast to PIC’s fixed segmentation target label point map, the in-context labeling training strategy introduces variability in the label point map for each point cloud. This difference poses unique challenges during evaluation, particularly when the target point cloud for the query is not visible. To address this issue, we leverage the label point bank $B_{j}$ utilized by the prompt $Q_{j}$ as the segmentation standard for the query.
During evaluation, for each point in the predicted point cloud, we compute its distance from each point in the label point set $B_{j}$. By selecting the label point with the minimum distance, we determine the segmentation result for each point. This process enables us to transform the dynamic regressed point cloud $\hat{P}_{i}$ into corresponding classification prediction while avoiding the fixed property from classification.
Loss Function. As the Chamfer Distance [chamferdistance] (CD) loss only measures the similarity between two point clouds, in our segmentation benchmark, we require precise regression of predictions for each point. Therefore, we incorporate Smooth-$\ell_{1}$ [fastrcnn] loss to assist the model in optimization. At the same time, our training purpose is to minimize the distance between the predicted point cloud and the target point cloud, so the CD loss in PIC is also used. The loss function of PIC++ can be formulated as:
5 Experiments
We first provide specific experimental details, including model parameters and training procedures, and information on other related methods (Sec. 5.1). Subsequently, we quantitatively and qualitatively evaluate the proposed PIC and PIC++ models on the ShapeNet In-Context Dataset (Sec. 5.2) and Multi-Entity In-Context Dataset (Sec. 5.3), comparing them with other methods. The focus lies in exploring their performance in multitasking and part segmentation scenarios. Finally, we conduct detailed ablation studies to examine the rationality of the model architecture design and selection (Sec. 5.4).
Table 3: Influence of data scale. We conduct experiments that include large-scale datasets Objaverse [objaverse] in the training set and analyze the impact of large-scale data on the model’s in-context learning ability. The blue arrow indicates the performance improvement brought by adding the Objaverse dataset.
(a) Experimental results of PIC on ShapeNet In-Context Datasets.
[…]
”-Cat” 4.3 5.3 14.1 78.95
4.3- 5.2↓0.1 10.7↓3.4 80.13↑1.18
”-Sep” 4.7 7.6 10.3 74.95
4.4↓0.3 6.2↓1.4 8.2↓2.1 81.96↑7.01
(b) Generalization results of PIC on ModelNet40 [modelnet40] on Registration. Note that ModelNet40 is not included in the training set. We report the CD Loss ($\times$1000).
Variants Obajverse L1 L2 L3 L4 L5 Average
”-Cat” 23.7 25.9 27.4 28.6 29.0 26.9
20.4↓3.3 22.3↓3.6 22.9↓4.5 21.6↓7.0 23.3↓5.7 22.1↓4.8
”-Sep” 17.6 20.1 18.1 16.4 15.3 20.3
15.3↓2.3 17.2↓2.9 17.1↓1.0 14.3↓2.1 12.8↓2.5 15.3↓5.0
5.1 Experimental Details
Implementation Details. We sample $N=1024$ points of each point cloud and divide it into $N_{C}=64$ point patches, each with $M=32$ neighborhood points. Besides, we set the mask ratio as $0.7$. We randomly select a prompt pair that performs the same task with the query point cloud from the training set. We consider $R_{rec}\in\{3\%,6\%,12\%,25\%,50\%\}$ and $R_{den}\in\{10\%,20\%,30\%,40\%,50\%\}$. We use an AdamW optimizer [adam] and cosine learning rate decay, with the initial learning rate as 0.001 and a weight decay of 0.05. For PIC++, we use a pre-trained checkpoint from PIC as model initialization. All models are trained for 300 epochs.
Baseline Methods and Training Details To evaluate the proposed framework, we evaluate the two variants (“-Sep” and “-Cat”) of PIC and PIC++ on the datasets mentioned in Sec. 3. We compare them with the following methods:
1) Task-Specific Models. We select three representative methods: PointNet [pointnet], DGCNN [dgcnn], PCT [pct], and ACT [act], and train them individually on each task or dataset. For the multitasking benchmark, we also design task-specific heads for each task.
2) Multitask Models. For a fair comparison, we develop multitask models based on common point cloud analysis methods. These models feature a shared backbone network and multiple task-specific heads designed to be capable of multitasking learning.